OpenStudy (anonymous):

A scale on a map shows that 2 inches = 25 miles. How many inches represents 60 miles?

OpenStudy (alexandervonhumboldt2):

x/60=2/25

OpenStudy (alexandervonhumboldt2):

x=2/25*60

OpenStudy (anonymous):

Wuuut. Thats what i got.

OpenStudy (anonymous):

Maybe im getting the second part of the question wrong :/

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

Ya, i was getting the other part of it wrong :/ Derp. I had 4.8